Reports 1-1 of 1 Clear search Modify search
Injection system (General activities)
chiummo, DeRossi, Gherardini - 23:35 Friday 07 April 2023 (59703) Print this report
Some BPC/PSTAB noise injections

Some noise injections have been performed on BPC and PSTAB degrees of freedom while at LOW_NOISE_1:

On VPM/ISYSnoise
#Driving matrix
BPC_TX_NOISE 0.0001 (on BPC_NOISE)

1364917605  07 Apr 2023 15:46:27 UTC  (07 Apr 2023 17:46:27 CEST)
Increase TX to 0.0020

1364918098  07 Apr 2023 15:54:40 UTC  (07 Apr 2023 17:54:40 CEST)
disable noise

there is coherence now and a clear increase of noise on DARM/Hrec starting from 100Hz (see first plot) but the noise is well visible also on other degrees of freedom, driving on BPC is not very well diagonalized.

BPC_TY_NOISE 0.0005 on BPC_NOISE
1364918386  07 Apr 2023 15:59:28 UTC  (07 Apr 2023 17:59:28 CEST)
enable colored noise
(no clear increase of coherence with either darm or hrec)

Increase TY to 0.0020
1364918564  07 Apr 2023 16:02:26 UTC  (07 Apr 2023 18:02:26 CEST)
300 sec data
no increase of noise on hrec darm


Increase TY to 0.01
1364919045  07 Apr 2023 16:10:27 UTC  (07 Apr 2023 18:10:27 CEST)
Disabled noise 16:16:24 (300 sec data)

Second plot. This time the increase of noise is more evident also at lower frequency, there is also a resonance excited around 800Hz - could be related to the piezo actuators.

BPC_X_NOISE 0.1 on BPC_NOISE
1364921543  07 Apr 2023 16:52:05 UTC  (07 Apr 2023 18:52:05 CEST)
(large coherence at 1KHz)
1364922219  07 Apr 2023 17:03:21 UTC  (07 Apr 2023 19:03:21 CEST)
disabled noise

Third plot. Noise increase evident only close to the resonance.

BPC_Y_NOISE 0.05 on BPC_NOISE
1364922484  07 Apr 2023 17:07:46 UTC  (07 Apr 2023 19:07:46 CEST)
enable
1364922799  07 Apr 2023 17:13:01 UTC  (07 Apr 2023 19:13:01 CEST)
disable

Fourth plot. Some noise increase in the 100Hz region.

 

Then we moved to PSTAB noise injections.

On ISYS Noise
Noise for PSTAB
ACL_SUM_CH [...] 10 PSTAB_colored 
1364924406  07 Apr 2023 17:39:48 UTC  (07 Apr 2023 19:39:48 CEST)
1364924758  07 Apr 2023 17:45:40 UTC  (07 Apr 2023 19:45:40 CEST)
disabled

    unlock for no clear reason
    1364924893  07 Apr 2023 17:47:55 UTC  (07 Apr 2023 19:47:55 CEST)

Then relocked at the first attempt.
Injection on fifth plot, low frequency.


Changed PSTAB_flt in ISYSnoise_filter (switched to 1KHz bandwidth)
ampl 0.1 on PSTAB colored
1364930774  07 Apr 2023 19:25:56 UTC  (07 Apr 2023 21:25:56 CEST)
1364931095  07 Apr 2023 19:31:17 UTC  (07 Apr 2023 21:31:17 CEST)
Noise disabled

Sixth plot was taken few seconds after switching off the noise, the level of the injected noise was instead comparable to the one used in the latest noise projections.

By mistake, also white noise was injected (seventh plot):

On ISYS Noise
Noise for PSTAB
ACL_SUM_CH [...] 0.005 PSTAB_white [...]
1364929392  07 Apr 2023 19:02:54 UTC  (07 Apr 2023 21:02:54 CEST)
1364929942  07 Apr 2023 19:12:04 UTC  (07 Apr 2023 21:12:04 CEST)
noise disabled 
(for PSTAB white noise there is no button enable/disable, reload config will directly apply the change)

 

I also tried to activate the permanent monitoring lines on BPC, without success.

After some rough evaluation, it seems that the BPC could be closer to DARM around 100Hz region. Lines moved to:
TX -> 97Hz
TY -> 89Hz
X -> 83Hz
Y -> 79Hz

Changes applied in ISYSnoise_filters
1364934495  07 Apr 2023 20:27:57 UTC  (07 Apr 2023 22:27:59 CEST)
Reload filters, tried to send the lines to the BPC corrections but something went wrong, will try to debug offline.

Interferometer left locked at LOW_NOISE_1.

Images attached to this report
Comments to this report:
flaminio - 10:25 Saturday 08 April 2023 (59707) Print this report

Very interesting measurements!

There is one thing to keep in mind. The BPC is measuring the beam jitter at the BPC quadrants on EIB. The beam jitter at the IMC input might be larger since the beam is travelling in air from the BPC quadrants to the IB tower input.

Maybe a reshaping of the IMC AA loops filters can reduce the noise at 100Hz.

This said, there is still one thing that I do not understand. The IMC AA WFS is measuring the misalignement between the noisy in air input beam and the silent suspended under vacuum IMC. Three of the four IMC AA WFS signals are correctly used to act on the BPC i.e. in the noisy beam. But the fourth acts on the IMC payload. Wouldn't be possible to use also the fourth to act on the last BPC degree of freedom i.e. BPC_Y? 

Given that the two MC mirror angular degrees of freedom can be controlled using a dithering like system that keeps the beam in the center of the MC mirror, we would then be left with IB_TX. Can this remain under local control?

If not, can a blending of the fourth IMC AA WFS between IB_TX (at the IB suspension resonant frequencies) and BPC_Y (at DC and at high frequency) be the solution?

capocasa, barsuglia - 13:02 Saturday 08 April 2023 (59709) Print this report

Fig.1 shows the comparison between the RIN transfer function (input beam -> B1_DC) measured yesterday (2023,04,07,19,26,00) and that measured last week (023,03,29,20,46,50).

The input RIN is computed as PSTAB_PDd_AC_MONIT/ (11*12*PSTAB_PDd_DC), while RIN_OUT is simply B1_DC/mean(B1_DC);

The TF has been compensated for the DARM closed loop (/virgoDev/ISC/matlab/DARM_loop_TF.m).

The coupling seems reduced by about a factor 2.

Fig. 2 shows the CMRF (DARM_SSFS_LINE_mag_100Hz) comparison in the two moments. The fluctuation is quite large but the mean values are respectively 2.7e-9 (old measurement) and 1.3e-9 (new measurement). This seems in agreement with the amplitude noise coupling reduction.

Images attached to this comment
mwas - 18:38 Saturday 08 April 2023 (59712) Print this report

The measured shape of the RIN transfer function is compatible with an offset compared to dark fringe of the BS.

Figure 1 shows a simple simulation in optickle of the transfer function from laser amplitude noise to B1 RIN (so there is a factor 2 compared to laser RIN to B1 RIN transfer function), with in red a perfect interferometer, and in blue an interferometer with a large BS offset (which make the power on B1 have roughly equal contribution from the DARM offset and from the MICH offset). The shape with the MICH offset above 100Hz looks very similar to what is measured. Compared to a perfect interferometer the impact of a MICH offset on RIN coupling is mainly above 100Hz.

FIgure 2 shows from the same simulation the transfer function from the laser phase (in radians) to B1 RIN. The impact of a MICH offset is mainly below 100Hz.

This suggest that one should use the coupling of PSTAB to DARM at high frequency (the 1.5kHz PSTAB line for example), to adjust the MICH offset. Given the large fluctuation we see at each lock, it would be good to have this a signal in loop adding a MICH offset once we are in LN1. The other tunings (TCS differential, Etalon, ...), could help the working point so that this position also corresponds to a minimum in SSFS coupling and a good 56MHz balancing on B1p, which currently is not the case.

Images attached to this comment
mwas - 21:54 Sunday 09 April 2023 (59719) Print this report

These are nice BPC injections, and I have updated the times in the noise budget to include them.

Figure 1 shows the result of the noise projection at a calm time across the frequency band, taking the transfer function where during the measurement the coherence with DARM is at least 30%.

Figure 2 is a zoom at low frequency. The bumps that are slighly below 20Hz, and which I had noticed also in MICH and SRCL, seem to be due to BPC TY. Note that it doesn't look present in h(t), presumably Hrec sees the coherence between DARM and MICH, and subtracts this way the BPC TY noise. It would be interesting to undertand better where this bump at ~18Hz is coming from.

Images attached to this comment
bertolini - 13:20 Monday 10 April 2023 (59722) Print this report

The bump mentioned by Michal is highly coherent with the ENV accelerometers on both LB and EIB. It's not completely clear whether the dominant contribution comes from. The reason is that both LB and EIB have structures around 17Hz; what i mean is that EIB is way quieter than LB along X but not so much along Z and it hosts the BPC sensors. As input for further investigation the second plot shows the EIB displacement noise in m/sqrt[Hz] units obtained by double integrating the ENV accelerometers signals.

Images attached to this comment
Tringali, Chiummo, Gosselin, Spinicelli - 23:52 Saturday 15 April 2023 (59807) Print this report

This afternoon, we investigated the bump around 18 Hz visible in DARM and its coherence with LB and EIB accelerometers. We have installed a temporary microphone ENV_LR_MIC in the laser room savaging the connection of the accelerometer ENV_LB_ACC_Z, Figure 1. We found the PCB accelerometer and microphone removed from the laser bench and left on the top of the plexiglass, Figure 2. We reinstalled the microphone, Figure 3. The accelerometer installation will be done on Tuesday morning (maintenance).
We switched off the air conditioning for a few minutes, from 15:56 to 16:02 UTC. We did not observe significant temperature variation inside the laser laboratory.
Figure 4 and Figure 5 (zoom) show the effect of air conditioning switched-off (blue curves) on the microphones (first row), accelerometers on the benches (second and third rows), Hrec and DARM (fifth row).  A broadband noise reduction is well visible on the environmental sensors while it seems to be a small reduction of Hrec and DARM around 18 Hz.
In Figure 6, we observe a high coherence between ENV_LR_MIC and Hrec, DARM when the air conditioning is on but it decreases abruptly during the switched-off (blue curves).
Regarding the coherence of the microphone in the laser room and the accelerometers on the benches, even in this case, we observe a reduction but some structures of coherence are still present when the air conditioning is switched off.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×